# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

### v2

@@ I INTRODUCTION
Although reinforcement learning-based models for robotic manipulation can learn to infer actions directly from image observations through interaction with the environment, they still heavily rely on carefully positioned static third-person or wrist-mounted cameras[1, 2, 3]. While these cameras reduce sample variance during interaction[3, 4], they often struggle with providing sufficient visual data in occluded or limited-visibility scenarios due to unadjustable viewpoints or fixed camera-gripper configuration.
However, humans can adjust their head and eye positions based on task goals. This helps them obtain advantageous visual information for action decision [5]. Inspired by this, recent active vision-action models enable agents to actively adjust their sensory viewpoints while interacting with the environment[3, 6, 7, 8], mimicking human capabilities.
Active vision-action models, like human perception, can be categorized as synchronous or asynchronous[10]. Synchronous models make decisions about visual and motor actions simultaneously based on current observations[7, 8, 11, 12]. However, the need for simultaneous coordination of sensor-motor actions places high demands on the model’s dynamic modeling capabilities and may limits its ability to handle tasks requiring large viewpoint shifts.
@@
We train our model with few-shot reinforcement learning[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories. To address the expanded observation-action space caused by the active viewpoint, we employ a viewpoint-centric voxel alignment technique that rotates the scene grid to align with the camera viewpoint. We also integrate a task-agnostic auxiliary reward to stabilize training, encouraging agents to interact effectively and select informative viewpoints for gripper action decisions. We evaluate our model on 8 RLBench[9] tasks with limited viewpoints, demonstrating its effectiveness in viewpoint selection and action inference based on task requirements. The main contributions of this paper are as follows:
@@
  • Comprehensive experiments on public RLBench tasks validate the effectiveness and robustness of our approach.
@@ II Related Work
Active Vision for Manipulation. Active vision aims to enable agents to actively observe the environment based on specific task goals. It has been widely applied in tasks such as indoor visual navigation[12, 15], object searching[16, 17] and 3D reconstruction[18]. In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7, 13, 14, 19] and for detecting fruit nodes in automated harvesting[20, 21, 22]. Recent studies have extended the use of active vision to more general robotic manipulation tasks. For instance, Cheng et al[11] proposed an active vision model with a target detector module to handle object-pushing task in occluded scenes. Lv et al[8] introduced a model-based sensing-aware framework for manipulating flexible objects. Matthew et al[4] employed a self-prediction approach to learn viewpoint adjustments while simultaneously developing motion skills for robotic arm joints. Similarly, Shang et al[6] utilized a self-supervised reward mechanism to jointly learn visual sensing and motor skills. Unlike these methods, our active vision-aciton model does not require a task-specific target detector and employs an RL-based dual agent asynchronous decision-making structure.
@@
Attention-driven Robotic Manipulation. The attention mechanism, widely applied in fields such as object detection and natural language processing, contributes to robotic manipulation by focusing on the most task-relevant scene information from raw image inputs, improving both training efficiency and model performance[30]. ARM [1] and C2FARM [2] use attention mechanisms to focus on key visual input regions, with ARM selecting noteworthy 2D regions via DQN, while C2FARM predicts regions of interest(ROIs) in 3D space. Liang et al. [31] pre-train a scene saliency attention encoder offline using human-annotated saliency maps and then train the reinforcement learning operation agent online based on the scene features output by this encoder. However, these methods largely rely on single or multiple fixed cameras from different viewpoints to capture images of the scene, lacking the capability to actively choose their viewpoints. Our model partially incorporates the attention mechanism from C2FARM and endows the model with the ability to predict the optimal task-related observation viewpoint, further mimicking the human’s active visual perception capability.
Figure 3: Overview of our model pipeline. The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optimal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal. The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
@@ III Problem Formulation
We divide one interaction time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy $\pi_{\text{v}}$ from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy $\pi_{\text{p}}$ from $t^{\prime}\sim t+1$. The asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},P^{\prime},r,\gamma>$, as shown in Fig.2. where $\mathcal{O}$ is the observation space, $\mathcal{A^{\text{c}}},\mathcal{A^{\text{g}}}$ is the NBV and NBP action space, respectively. At time $t$, the NBV policy $\pi_{\text{v}}$ selects a camera action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, resulting in a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ via transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$. The NBP policy $\pi_{\text{p}}$ then selects a gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$, with the next scene observation $\mathbf{o}_{t^{\prime}}$ obtained through $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$. The environment provides a reward $r_{t+1}\in\mathbb{R}$. Our goal is to jointly learn $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{p}}$ to maximize the expected sensor-motor reward $\mathbb{E}_{\pi_{\text{v}},\pi_{\text{p}}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
Observation and Action Space. We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid derived from the RGB-D image, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state. Specifically, the RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, and voxelized via $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center. The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and an occupancy flag for each voxel. The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
Following common practice[14, 19], the camera’s movement is limited to the upper hemisphere at a fixed radius $r^{\text{v}}=1.2\text{m}$ above the table, with the optical axis directed at the sphere center, illustrated in Fig.8(b). The viewpoint is described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$. The NBV policy action is $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene. The NBP policy action $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}$ consists of the discrete 6-DoF gripper pose index $\mathbf{k}^{\text{g}}$ and the goal distance $d$ between the gripper tips.
@@ IV-A Model Structure
We illustrates the basic pipline of our model in Fig3. Once the alignment process is completed(Section IV-B), the observation $\mathbf{o}_{t}$ is obtained. The DQN-based NBV policy network, parameterized by $\Theta$, then uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1). Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, which represents the viewpoint most likely to enhance task success. The camera is subsequently adjusted to the goal viewpoint $\mathbf{v}^{*}$, derived from the selected $\mathbf{k}^{\text{v}}$.
$\mathbf{a}_{t}^{\text{c}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}})$ (1)
Based on the ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the image from $\mathbf{v}^{*}$ is cropped and aligned to generate the 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$ for the ROI. The NBP policy network, also DQN-based and parameterized by $\Phi$, infers the gripper closure $d$ and pose index $\mathbf{k}^{\text{g}}$ based on $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$. Finally, a complete sensor-motor interaction step is achieved after gripper interaction and rewarded by $r_{t+1}$.
$\mathbf{a}_{t^{\prime}}^{\text{g}}=\arg\underset{\mathbf{k}^{\text{g}},d}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},d)$ (2)
We denote a complete interaction transition as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$, where $\text{T}_{t}^{\text{v}}=(\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}})$ represents the NBV policy’s sub-transition, and $\text{T}_{t}^{\text{p}}=(\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1})$ represents the NBP policy’s sub-transition. After each training episode, transitions are stored in the replay buffer along with demonstration data. Both policies are updated using the same task reward $r_{t+1}$, thereby achieving sensor-motor joint learning. Each policy employs a separate 3D-CNN based U-Net[32] as the image encoder, followed by a three-layer fully connected network(FCN) for action decoding.
[…]
@@ IV-B Viewpoint-centric Voxel Alignment
$\mathbf{o}^{\text{v}}=\mathbf{V}\big{(}(^{\text{L}}{\text{T}}_{\text{W}}\times\mathbf{o}^{\text{pc}}),\mathbf{e},\mathbf{c}\big{)}$ (3)
@@
$\mathbf{v}^{*},\mathbf{f}^{*},\mathbf{g}^{*}=^{\text{W}}{\text{T}}_{\text{L}}\times(\mathbf{v}^{*}_{\text{L}},\mathbf{f}^{*}_{\text{L}},\mathbf{g}^{*}_{\text{L}})$ (4)

### v3

@@ Abstract
Index Terms:
Manipulation Planning; Reinforcement Learning; Perception for Grasping and Manipulation.
@@ I Introduction
Although reinforcement learning-based models for robotic manipulation can learn to infer actions directly from image observations through interaction with the environment, they still heavily rely on carefully positioned static third-person or wrist-mounted cameras[1, 2, 3]. While these setups reduce sample variance during interaction[3, 4], they often struggle with providing sufficient visual data in occluded or limited-visibility scenarios due to unadjustable viewpoints or fixed camera-gripper configuration.
Alternatively, humans can adjust their head and eye positions according to task goals, which helps them obtain advantageous visual information for action decision-making [5]. Inspired by this, recent active vision-action models allow agents to actively adjust their sensory viewpoints while interacting with the environment [3, 6, 7, 8]. These models, which mimic human capabilities, can be categorized as either synchronous or asynchronous [9].
Synchronous models make decisions about visual and motor actions simultaneously based on current observations[7, 8, 10, 11]. However, the need for simultaneous coordination of sensor-motor actions places high demands on the model’s dynamic modeling capabilities and may limits its ability to handle tasks requiring large viewpoint shifts.
@@
We train our model with few-shot reinforcement learning[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories. To address the expanded observation-action space caused by the active viewpoint, we employ a viewpoint-centric voxel alignment technique that rotates the scene voxel grid to align with the adjusted viewpoint. We also integrate a task-agnostic auxiliary reward to stabilize training, encouraging agents to interact effectively and select informative viewpoints for gripper action decisions. We evaluate our model on 8 RLBench[12] tasks with limited viewpoints, demonstrating its effectiveness in viewpoint selection and action inference based on task requirements. The main contributions of this paper are as follows:
@@
  • Comprehensive experiments on 8 public RLBench tasks validate the effectiveness and robustness of our approach.
@@ II Related Work
Active Vision for Manipulation. Active vision [15] aims to enable agents to actively observe their environment based on specific task goals, such as visual navigation[11], object recognition[16, 17], and 3D reconstruction [18, 19]. In robotic manipulation, active vision is primarily applied to tasks such as object grasping [7, 13, 14, 20] and fruit harvesting [21, 22]. For more general manipulation tasks, Cheng et al. [10] trained an RL-based viewpoint policy to detect occluded objects and perform a pushing task. Lv et al. [8] trained a model-based active sensing-action model for flexible object manipulation. Chuang et al. [23] used imitation learning to train an active viewpoint strategy that provides low-occlusion observations for peg insertion tasks. Shang et al. [6] developed a self-supervised reward mechanism to jointly learn active visual sensing and motor skills for door opening and table wiping tasks. Dengler et al. [24] trained an active viewpoint policy to learn how to push objects in confined spaces. Our work is most similar to [10], but unlike their model, we do not rely on a task-specific object detector. Furthermore, our active vision-action model uses an RL-based dual-agent asynchronous decision-making structure, allowing the agent to observe the environment first and then interact with it, similar to human behavior.
@@
Attention-driven Robotic Manipulation. The attention mechanism, widely applied in fields such as object detection and natural language processing, contributes to robotic manipulation by focusing on the most task-relevant scene information from raw image inputs[32], improving both training efficiency and model performance. Gualtieri et al. [33] generate grasp actions by applying attention to point cloud images to select noteworthy local regions and performing fine perception on these regions. ARM [1] and C2FARM [2] use attention mechanisms to focus on key visual input regions, with ARM selecting noteworthy regions from 2D inputs, while C2FARM predicts 3D regions of interest (ROI) from voxelized 3D images. Liang et al. [34] pre-train a scene saliency attention encoder offline using human-annotated saliency maps and then train the reinforcement learning operation agent online based on the scene features output by this encoder. However, these methods largely rely on single or multiple fixed cameras from different viewpoints to capture images of the scene, lacking the capability to actively choose their viewpoints. Our model incorporates the attention mechanism from C2FARM to select noteworthy local 3D regions from the scene but further enhances our model with the ability to predict the optimal observation angle for these regions, further mimicking human active visual perception capabilities.
Figure 3: Illustration of model pipeline and action spaces. The NBV policy infers the 3D ROI position $\mathbf{r}^{*}$ based on the global scene observation and predict the optimal viewpoint $\mathbf{v}^{*}$ for the ROI observation according to the given task goal. The NBP agent determines the gripper actions based on the ROI observation from $\mathbf{v}^{*}$.
@@ III Problem Formulation
We divide one interaction time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy $\pi_{\text{v}}$ from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy $\pi_{\text{p}}$ from $t^{\prime}\sim t+1$. The asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{S},\mathcal{O},\mathcal{A^{\text{v}}},P,\mathcal{A^{\text{p}}},P^{\prime},r,\gamma>$, as shown in Fig.2, where $\mathcal{S}$ is the environment states, $\mathcal{O}$ is the observation space, $\mathcal{A^{\text{v}}},\mathcal{A^{\text{p}}}$ is the NBV and NBP action space, respectively. At time $t$, the NBV policy $\pi_{\text{v}}$ selects a viewpoint action $\mathbf{a}_{t}^{\text{v}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{v}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, resulting in a new observation $\mathbf{o}_{t^{\prime}}\sim P(\cdot|\mathbf{s}_{t},\mathbf{a}_{t}^{\text{v}})$. The NBP policy $\pi_{\text{p}}$ then selects a gripper action $\mathbf{a}_{t^{\prime}}^{\text{p}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{p}}}$, with the next scene observation $\mathbf{o}_{t+1}\sim P^{\prime}(\cdot|\mathbf{s}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{p}})$ and receives a reward $r_{t+1}\in\mathbb{R}$. Our goal is to jointly learn $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{p}}$ to maximize the expected sensor-motor reward $\mathbb{E}_{\pi_{\text{v}},\pi_{\text{p}}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\cdot r(\mathbf{s}_{t},\mathbf{a}^{\text{v}}_{t},\mathbf{a}^{\text{p}}_{t^{\prime}},\mathbf{s}_{t+1})\Big{]}$.
Observation and Action Space. The observation spaces $\mathcal{O}=\{\mathcal{O}^{\text{i}},\mathcal{O}^{\text{s}}\}$, where $\mathbf{o}^{\text{i}}\in\mathcal{O}^{\text{i}}$ represents the 3D voxelized scene image from the RGB-D data, and $\mathbf{o}^{\text{s}}\in\mathcal{O}^{\text{s}}$ is the proprioceptive state. Specifically, the 3D voxel grid $\mathbf{o}^{\text{i}}\in\mathbb{R}^{n^{2}\times(3+M+1)}$ contains the voxel coordinates, $M$-dimensional features, and one occupancy flag for each voxel, where $n=16$ and $M=11$. The proprioceptive state $\mathbf{o}^{\text{s}}=\{\mathbf{v},\mathbf{g},d\}$ comprises the viewpoint spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$, the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$ and the gripper finger tips distance $d\in\mathbb{R}$. The camera’s movement is limited to the upper hemisphere at a fixed radius $r^{\text{v}}=1.2\text{m}$ above the table, illustrated in Fig.3, with the optical axis directed at the sphere center. The NBV action $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{r}}\}$, where $\mathbf{k}^{\text{v}}\in\mathbb{Z}^{2}$ is the discrete viewpoint index for $\theta^{\text{v}}\in[10^{\circ},60^{\circ}]$ and $\phi^{\text{v}}\in[-135^{\circ},135^{\circ}]$ with a resolution of $5^{\circ}$. $\mathbf{k}^{\text{r}}\in\mathbb{Z}^{3}$ denotes the position index of the 3D ROI that requires attention within the workspace. The $1\text{m}^{3}$ workspace is discretized into $16^{3}$ uniform grids, resulting in each dimension of $\mathbf{k}^{\text{r}}$ having an index range of $[0,15]$. The NBP action $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},\mathbf{k}^{\text{c}}\}$ consists of the discrete gripper pose index $\mathbf{k}^{\text{g}}\in\mathbb{Z}^{6}$ and desired gripper closure state $\mathbf{k}^{\text{c}}\in[0,1]$. The gripper position is constrained within the $0.3\text{m}^{3}$ ROI, which is further discretized into $16^{3}$ grids. The gripper’s rotation is represented by Euler angles on each axis, discretized in the range $[0^{\circ},360^{\circ}]$ with a resolution of $5^{\circ}$. It is important to note that while both NBV and NBP agents share the same observation format, NBV observes the entire scene, whereas NBP focuses only on the 3D ROI image, achieving an attention effect [2].
@@ IV-A Model Structure
Both the NBV and NBP policies are implemented using a DQN network with a 3D U-Net backbone[35], parameterized as $\Theta$ and $\Phi$, respectively. The NBV agent infers the center position index $\mathbf{k}^{\text{r}}$ of the ROI in the scene and the optimal viewpoint index $\mathbf{k}^{\text{v}}$ for the ROI region based on the preprocessed aligned observation $\mathbf{o}_{t}$, as shown in Eq.(1). It then interacts with the environment and moves to move the viewpoint to the corresponding goal position $\mathbf{v}^{*}$.
$\mathbf{a}_{t}^{\text{v}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{r}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{r}})$ (1)
The depth image data captured at time $t^{\prime}$ is cropped and voxelized, yielding the voxel grid observation $\mathbf{o}^{\text{i}}_{t^{\prime}}$ for the ROI region under the new viewpoint. The NBP agent then infers the gripper pose index $\mathbf{k}^{\text{g}}$ and closure action $\mathbf{k}^{\text{c}}$ based on observation $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{i}}_{t^{\prime}},\mathbf{o}^{\text{s}}_{t^{\prime}}\}$.
$\mathbf{a}_{t^{\prime}}^{\text{p}}=\arg\underset{\mathbf{k}^{\text{g}},\mathbf{k}^{\text{c}}}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},\mathbf{k}^{\text{c}})$ (2)
Finally, the arm trajectory is planned using an off-the-shelf trajectory planner and executed through multiple control steps to move the gripper to the goal position $\mathbf{g}^{*}$, completing one sensor-motor interaction and receiving the reward $r_{t+1}$. The transition of the global agent denote as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$ , where $\text{T}_{t}^{\text{v}}=(\mathbf{o}_{t},\mathbf{a}_{t}^{\text{v}},\mathbf{o}_{t^{\prime}})$ is the local transition of the NBV agent, and $\text{T}_{t}^{\text{p}}=(\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{p}},\mathbf{o}_{t+1})$ is the NBP agent’s local transition. The explored transitions are stored in the replay buffer along with demonstration data during training.
@@ IV-B Viewpoint-centric Voxel Alignment
$\mathbf{o}^{\text{i}}=\mathbf{vox}(^{\text{L}}{\text{T}}_{\text{W}}\times\mathbf{o}^{\text{pc}})$ (3)
@@
[…]
$\mathbf{v},\mathbf{r},\mathbf{g}=^{\text{W}}{\text{T}}_{\text{L}}\times(\mathbf{v}_{\text{L}}^{*},\mathbf{r}_{\text{L}}^{*},\mathbf{g}_{\text{L}}^{*})$ (4)